OpenStudy (anonymous):

x^2-16x+7=0 @StudyGurl14

OpenStudy (studygurl14):

The quadratic formula: Where you have \(\large a^2+bx+c=0\) \(\large x=\Large\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (studygurl14):

If you have \(\large x^2-16x+7=0\) What do a, b, and c equal? @kat_lee17

OpenStudy (anonymous):

a= 1 b= -16 c=7 thanks

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

correct. plug those values into the formula

OpenStudy (anonymous):